12.0 moles of gas are in a 3.00L tank at 23.1?C . Calculate the difference in pressure between methane and an ideal gas under these conditions. The van der Waals constants for methane are a=2.300L2?atm/mol2 and b=0.0430 L/mol.
Answer #1

Given, moles = n = 12.0 mol

Volume = v = 3.00 L

Temperature = T = 23.1 degree C = 296.1 K

Ideal gas equation => PV = nRT

P = nRT /V

P = (12.0) ( 0.0821) (296.1) / 3.00

P = 97.24 atm.

--------------------

Given,

Vanderwaal constants

a=2.300 L2 atm/mol2 and

b=0.0430 L/mol.

moles = n = 12.0 mol

Volume = v = 3.00 L

Temperature = T = 23.1 degree C = 296.1 K

[P + (n^2a/V^2)] [V-nb] = nRT

[P + (144 x 2.300) / 9] [ 3 - 0.516] = 12 x 0.0821 x 296.1

[P + 36.8] - 2.484 = 291.7

P = 257.40 atm

--------------------

Difference = P vanderwaal - P ideal

= 257.40 - 97.24

= 160.16 atm